Inverse Laplace Transforms

Find L^-1 {F(s)} when F(s) is given by


5. 1/(s+1)(s+2)(s^2+2s+10)


1
Expert's answer
2022-05-29T16:07:55-0400

5.


"\dfrac{1}{(s+1)(s+2)(s^2+2s+10)}=\dfrac{A}{s+1}+\dfrac{B}{s+2}"




"+\dfrac{Cs+D}{s^2+2s+10}=\dfrac{A(s+2)(s^2+2s+10)}{(s+1)(s+2)(s^2+2s+10)}"


"+\dfrac{B(s+1)(s^2+2s+10)}{(s+1)(s+2)(s^2+2s+10)}"

"+\dfrac{(Cs+D)(s+1)(s+2)}{(s+1)(s+2)(s^2+2s+10)}"

"s=-1:A(1-2+10)=1"

"s=-2:-B(4-4+10)=1"

"s=0: 20A+10B+2D=1"


"s=1:39A+26B+6C+6D=1"

"A=\dfrac{1}{9}, B=-\dfrac{1}{10}, D=-\dfrac{1}{9}, C=-\dfrac{1}{90}"


"L^{-1}(\dfrac{1}{(s+1)(s+2)(s^2+2s+10)})"

"=\dfrac{1}{9}L^{-1}(\dfrac{1}{s+1})-\dfrac{1}{10}L^{-1}(\dfrac{1}{s+2})"

"-\dfrac{1}{90}L^{-1}(\dfrac{s+1}{(s+1)^2+3^2})-\dfrac{1}{10}L^{-1}(\dfrac{1}{(s+1)^2+3^2})"



"=\dfrac{1}{9}e^{-t}-\dfrac{1}{10}e^{-2t}-\dfrac{1}{90}e^{-t}\cos(3t)-\dfrac{1}{30}e^{-t}\sin (3t)"